For the following investment, find the total number of compounding periods (n) and the interest rate per period (i) that you would substitute into the future value or present value formula . (Do not round.)

Time Annual Rate Compounded Rate per period (in decimal form) Compounding periods
8 years 4.5% Monthly                   
Answer #1

Answer: Rate per period=0.0375 and compounding period=96

Explanation: Using the general formula for compounding interest.

A=P(1+\frac{r}{n})^{nt}

where, nt is the total number of compounding periods and  \frac{r}{n} is the rate per period, where, r=4.5% is the interest rate (yearly).

For monthly compounding, n=12

So, for t=8 years, total number of compounding period=n*t=12*8=96.

and rate per period=r/n=4.5%/12=4.5/(12*100)=0.00375